OpenStudy (anonymous):

Find the value of 1-2+3-4+5-6+...-1000000.

OpenStudy (anonymous):

Sum of odd numbers 1,3,....., 999999 - Sum of even numbers 2,4,6,.....,1000000 Use formula from Arithmetic series

OpenStudy (anonymous):

SUM OF ODD NUMBERS: If N = number of odd terms then sum = (N)^2 SUM OF EVEN NUMBERS: Formula: N(N+1) How to Find N = (First Even + Last Even)/2 - 1

OpenStudy (anonymous):

I know the answer.

OpenStudy (angela210793):

u know how to solve it?"

OpenStudy (anonymous):

yes. Just wanted to see how people would tackle it, so that I give a few medals to people.

OpenStudy (angela210793):

ohhhhh :)

OpenStudy (anonymous):

Another way of doing this is to notice that 1-2 = -1, 1-2+3 = 2, 1-2+3-4 = -2, 1-2+3-4+5 = 3, 1-2+3-4+5-6 = -3, etc. So continuing until -1000000, answer should be -100000/2 = -500000.

OpenStudy (anonymous):

\[\sum_{n=1}^{1000000}\left( n \right)\left( -1 \right)^{(n+1)}\] you can also write the summation in terms of the series expansion above!

OpenStudy (angela210793):

tht's wht i was thinking Xactxx